A boat traveled 210 miles downstream and back. The trip dwonstream took 10hours. The trip back took 70 hours. What is the speed of the boat in still water? What is the speed of the current? Explain.
Answer by ankor@dixie-net.com(22740)

let s = the boat speed in still water
let c = speed of the current
then
(s+c) = effective speed down stream
and
(s-c) = effective speed upstream
:
A boat traveled 210 miles downstream and back.
The trip dwonstream took 10hours.
The trip back took 70 hours.
What is the speed of the boat in still water?
What is the speed of the current? Explain.
:
Write a distance equation for each way. dist = time * speed
10(s+c) = 210
70(s-c) = 210
simplify both equation, divide the 1st by 10, the 2nd by 70
s + c = 21
s - c = 3
---------------addition eliminates c, find s
2s + 0 = 24
s = 24/2
s = 12 mph in still water
Find c
12 + c = 21
c = 21 - 12
c = 9 mph is the speed of the current
;
:
:
Check solutions in the 2nd original equation
70(12 - 9) = 210
70(3) = 210 mi